# Random Turán Problems for Hypergraph Expansions

### v1

@@ Abstract
The random Turán number $\mathrm{ex}(G_{n,p}^{r},F)$ is the maximum number of edges in an $F$-free subgraph of the random $r$-uniform hypergraph $G_{n,p}^{r}$. We prove general results which (informally) shows that if $F$ is an $r_{0}$-graph, then upper bounds for $\mathrm{ex}(G_{n,p}^{r_{0}},F)$ can be lifted into upper bounds for $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ where $F^{(r)}$ is the $r$-uniform expansion of $F$, i.e. the hypergraph obtained from $F$ by inserting $r-r_{0}$ distinct vertices into each edge of $F$. These results unify and generalize most known upper bounds for random Turán numbers of degenerate hypergraphs of uniformity at least 3, and also provide new tight bounds for the random Turán numbers of expansions of theta graphs and complete bipartite graphs.
@@ 1 Introduction
Given an $r$-uniform hypergraph $F$ (or $r$-graph for short), we recall that the Turán number $\mathrm{ex}(n,F)$ is the maximum number of edges that an $n$-vertex $F$-free $r$-graph can have. Let $G_{n,p}^{r}$ denote the random $r$-graph on $n$ vertices obtained by including each possible edge independently and with probability $p$, and when $r=2$ we simply write $G_{n,p}$ instead of $G_{n,p}^{2}$. We define the random Turán number $\mathrm{ex}(G_{n,p}^{r},F)$ to be the maximum number of edges in an $F$-free subgraph of $G_{n,p}^{r}$. Note that when $p=1$ we have $\mathrm{ex}(G_{n,1}^{r},F)=\mathrm{ex}(n,F)$, so the random Turán number can be viewed as a probabilistic analog of the classical Turán number.
@@
The asymptotics for $\mathrm{ex}(G_{n,p}^{r},F)$ are essentially known if $F$ is not an $r$-partite $r$-graph due to independent breakthrough work of Conlon and Gowers [3] and of Schacht [21], and because of this we focus only on the degenerate case when $F$ is $r$-partite. This degenerate case seems to be very difficult even in the graph setting $r=2$ where effective upper bounds are known only for even cycles [10, 5], complete bipartite graphs [10], and theta graphs [8]; and effective lower bounds are known only for powers of rooted trees [22]. All of these bounds agree with a recent conjecture of McKinely and Spiro [8], which gives a general prediction for how the function $\mathrm{ex}(G_{n,p},F)$ should behave for all bipartite $F$.
@@
For $r$-graphs with $r\geq 3$, the only $F$ for which effective bounds are known are complete $r$-partite $r$-graphs [23], loose even cycles [13, 15], Berge cycles [15, 23], expansions of hypergraph cliques and subgraphs of tight trees [18, 14], and non-Sidorenko hypergraphs [17]. We note that part of the difficulty for hypergraphs lies in the fact that no analog of the McKinley-Spiro conjecture mentioned above is known to hold for $r$-graphs, which is closely related to the lack of a hypergraph analog for Sidorenko’s conjecture [17].
The primary aim of this paper is to expand on the current knowledge of random Turán numbers for hypergraphs by studying the following wide class of examples.
@@
Hyeprgraph expansions are also well-studied (though not in a systematic way) in the context of random Turán numbers, see for example [13, 18, 14, 15] mentioned above. This is in large part because, as was initially observed by Mubayi and Yepremyan [13], one can sometimes [^3] Provided these bounds come from a balanced supersaturation result; see Section 1.2 for more. take known bounds for the random Turán problem of $F$ and “lift” these to bounds for the random Turán problem of expansions $F^{(r)}$.
Somewhat more precisely, given an $r$-graph $H$, we say that a set of vertices $S$ with $|S|=r_{0}$ is an $r_{0}$-shadow of $H$ if there exists an edge of $H$ containing the set $S$. Mubayi and Yepremyan thus obtained their upper bounds for the random Turán number of loose even cycles by (roughly speaking) looking at the 2-shadows of $G^{r}_{n,p}$ and then applying known random Turán bounds for graph even cycles due to Morris and Saxton [10] as a black box. A similar (though ultimately inequivalent) approach was used by Nie [15] by inductively looking at the $(r-1)$-shadows of $G^{r}_{n,p}$ rather than the 2-shadows. In total, these two approaches combine to give the following bounds, where here we say that a sequence of events $A_{n}$ holds asymptotically almost surely or a.a.s. for short if $\Pr[A_{n}]\to 1$ as $n\to\infty$, and we write $f(n)\ll g(n)$ if $f(n)/g(n)\to 0$ as $n\to\infty$.
@@
We emphasize here the peculiar situation that tight bounds for loose cycles are not known when $r=3$ despite the tight bounds known for $r=2$ [10] and for $r\geq 4$ due to 1.1.
1.1 Main Results: Random Turán Problems
The aim of this paper is to give a systematic generalization of the approaches of Mubayi-Yepremyan and Nie for loose even cycles mentioned above in order to prove a series of “Lifting Theorems” which informally allows one to lift bounds for the random Turán problem from hypergraphs $F$ into bounds for expansions $F^{(r)}$. Our Lifting Theorems are technical to state, so we postpone their formal statements until the next subsection and instead discuss some of their applications here.
Our Lifting Theorems together with known results on the random Turán problem for graph even cycles $C_{2\ell}$ can be used to prove 1.1, and more generally, these theorems together with known results can be used to recover every previously known upper bound for random Turán problems of expansions [13, 18, 14, 15]. These Lifting Theorems can also be used to obtain many new results as well. For example, while tight bounds for $C_{4}^{(3)}=K_{2,2}^{(3)}$ remain unknown, we can prove tight bounds for all other $K_{2,t}$ expansions.
@@ 1.1 Main Results: Random Turán Problems
If $r,t\geq 2$ are integers such that either $r,t\geq 3$ or $r\geq 4$, then for all $0<p=p(n)\leq 1$ we have a.a.s.
@@
More generally, our results extend to all $K_{s,t}$ at the cost of a suboptimal bound on $r$.
@@
$\mathrm{ex}(G^{r}_{n,p},K_{s,t}^{(r)})=\left\{\begin{aligned} &\max\{\Theta(pn^{r-1}),n^{2+\frac{s+t-2}{st-1}}(\log n)^{\Theta(1)}\},~{}~{}&\text{if}~{}p\gg n^{-r+2+\frac{s+t-2}{st-1}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-r+2+\frac{s+t-2}{st-1}}.\\ \end{aligned}\right.$
It is possible that these tight bounds for $\mathrm{ex}(G_{n,p}^{r},K_{s,t}^{(r)})$ continue to hold for any $r$ such that (${r\choose 2}\geq s$ as this is precisely the range where $\mathrm{ex}(n,K_{s,t}^{(r)})=\Theta(n^{r-1})$ for $t$ sufficiently large [7, Theorem 1.4], but 1.3 is the best bound for $r$ that can be obtained with our Lifting Theorems and the known (tight) random Turán bounds for the graph $K_{s,t}$ [10]. Nevertheless, it is possible to go beyond 1.3 by applying our Lifting Theorems to the hypergraphs $K_{s,t}^{(3)}$ instead of the graphs $K_{s,t}$. We will do precisely this in forthcoming work [16] where we use our Lifting Theorems together with a somewhat involved supersaturation argument to obtain the tight bounds of 1.3 for $r$ at least roughly $32s/$ 3 for $s\geq 3$.
Finally, our Lifting Theorems gives bounds for expansions of theta graphs. Here the theta graph $\theta_{a,b}$ is defined to be the graph consisting of $a$ internally disjoint paths of length $b$ between a common pair of vertices $u,v$. For example, $\theta_{2,b}=C_{2b}$ and $\theta_{a,2}=K_{2,a}$. Our Lifting Theorems gives good bounds on $\theta_{a,b}^{(r)}$ whenever effective bounds for $\theta_{a,b}$ are known and $a\geq 3$ (see 5.2), and in particular, recent results of McKinley and Spiro [8] give the following.
@@
1.2 Main Results: Balanced Supersaturation
There are two key ingredients in proving upper bounds for the random Turán problem. The first is hypergraph containers, which is a general method established independently by Balogh, Morris, and Samotij [1] and by Saxton and Thomason [20] for upper bounding the number of independent sets in hypergraphs. The second is balanced supersaturation, which is a concept originally introduced by Morris and Saxton [10].
Informally, a balanced supersaturation result for an $r$-graph $F$ says that if an $r$-graph $H$ has many edges, then one can find a collection $\mathcal{H}$ of copies of $F$ in $H$ which is “spread out” in the sense that no set of edges in $H$ lies in too many copies of $F$ in $\mathcal{H}$. More precisely we have the following definition, where here for a hypergraph $\mathcal{H}$ and $S\subseteq V(\mathcal{H})$ we define the degree $\deg_{\mathcal{H}}(S)$ to be the number of edges of $\mathcal{H}$ containing $S$, and we define $\Delta_{i}(\mathcal{H})=\max\{\deg_{\mathcal{H}}(S):S\subseteq V(\mathcal{H}),\ |S|=i\}$.

### v2

@@ Abstract
Given an $r_{0}$-uniform hypergraph $F$, we define its $r$-uniform expansion $F^{(r)}$ to be the hypergraph obtained from $F$ by inserting $r-r_{0}$ distinct vertices into each edge of $F$, and we define $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ to be the largest $F^{(r)}$-free subgraph of the random hypergraph $G_{n,p}^{r}$.
We initiate the first systematic study of $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ for general hypergraphs $F$. Our main result essentially resolves this problem for large $r$ by showing that $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ goes through three predictable phases whenever $F$ is Sidorenko and $r$ is sufficiently large, with the behavior of $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ being provably more complex whenever $F$ has no Sidorenko expansion. Moreover, our methods unify and generalize almost all previously known results for the random Turán problem for degenerate hypergraphs of uniformity at least 3.
@@ 1 Introduction
This paper involves the intersection of three different streams of research in extremal combinatorics: random Turán problems, Sidorenko hypergraphs, and hypergraph expansions. Throughout this paper we will work with $r$-uniform hypergraphs, or $r$-graphs for short. Given an $r$-graph $F$, we let $v(F)$ denote the number of vertices of $F$ and $|F|$ the number of edges of $F$; and if $|F|\geq 2$, then we define its $r$-density $d_{r}(F)$ by
$d_{r}(F):=\max_{F^{\prime}\subseteq F,\ |F^{\prime}|\geq 2}\frac{|F^{\prime}|-1}{v(F^{\prime})-r}.$
1.1 Random Turán Problems
@@
The Turán number $\mathrm{ex}(n,F)$ of an $r$-graph $F$ is defined to be the maximum number of edges that an $n$-vertex $F$-free $r$-graph can have. Let $G_{n,p}^{r}$ denote the random $r$-graph on $n$ vertices obtained by including each possible edge independently and with probability $p$, and when $r=2$ we simply write $G_{n,p}$ instead of $G_{n,p}^{2}$. We define the random Turán number $\mathrm{ex}(G_{n,p}^{r},F)$ to be the maximum number of edges in an $F$-free subgraph of $G_{n,p}^{r}$. Note that when $p=1$ we have $\mathrm{ex}(G_{n,1}^{r},F)=\mathrm{ex}(n,F)$, so the random Turán number can be viewed as a probabilistic analog of the classical Turán number.
@@
The asymptotics for $\mathrm{ex}(G_{n,p}^{r},F)$ are essentially known if $F$ is not an $r$-partite $r$-graph due to independent breakthrough work of Conlon and Gowers [4] and of Schacht [33], and because of this, we focus only on the degenerate case when $F$ is $r$-partite. This degenerate case seems to be very difficult, even in the graph setting $r=2$ where tight bounds are known only for even cycles [22, 13], complete bipartite graphs [22], and theta graphs [20]. All of the tight bounds from these examples agree with the following conjecture of McKinley and Spiro [20], which gives a general prediction for how the function $\mathrm{ex}(G_{n,p},F)$ should behave for all bipartite $F$. To state this conjecture, we say that a sequence of events $A_{n}$ holds asymptotically almost surely or a.a.s. for short if $\Pr[A_{n}]\to 1$ as $n\to\infty$, and we write $f(n)\ll g(n)$ if $f(n)/g(n)\to 0$ as $n\to\infty$.
Conjecture 1 ([20]).
If $F$ is a graph with $\mathrm{ex}(n,F)=\Theta(n^{\alpha})$ for some $\alpha\in(1,2]$, then a.a.s.
$\mathrm{ex}(G_{n,p},F)=\begin{cases}\max\{\Theta(p^{\alpha-1}n^{\alpha}),n^{2-1/d_{2}(F)}(\log n)^{\Theta(1)}\}&p\gg n^{-1/d_{2}(F)},\\ (1+o(1))p{n\choose 2}&n^{-2}\ll p\ll n^{-1/d_{2}(F)}.\end{cases}$
In particular, this conjecture predicts that $\mathrm{ex}(G_{n,p},F)$ should always have three ranges of behavior (i.e. it should be roughly equal to either $p^{\alpha-1}n^{\alpha},\ n^{2-1/d_{2}(F)}$, or $p{n\choose 2}$ depending on the value of $p$), and moreover, it predicts that one of these ranges will be a “flat middle range,” i.e.  a range where $\mathrm{ex}(G_{n,p},F)=n^{2-1/d_{2}(F)}(\log n)^{\Theta(1)}$ is essentially independent of $p$ for a sizable range of $p$ close to $n^{-1/d_{2}(F)}$.
Very few results are known regarding 1 for general graphs. In particular, the only upper bounds known for $\mathrm{ex}(G_{n,p},F)$ for arbitrary graphs $F$ are due to Jiang and Longbrake [13], with these general bounds matching 1 when $F$ is an even cycle. However, it can be shown that in all cases other than $F=C_{2\ell}$, the bounds of [13] are weaker than what is predicted by 1, so more work is needed to prove the bounds of 1 in every other case.
@@
Turning now from graphs to hypergraphs, we note that not much is known about the random Turán problem for $r$-graphs with $r\geq 3$. Indeed, the only such $F$ for which effective (though typically non-tight) bounds are known for are complete $r$-partite $r$-graphs [36], loose even cycles [25, 27], Berge cycles [27, 36], expansions of hypergraph cliques and subgraphs of tight trees [30, 26], and non-Sidorenko hypergraphs [29].
1.2 Sidorenko Hypergraphs
Part of the difficulty in proving results about random Turán numbers for hypergraphs is that no analog of 1 is known to hold for $r$-graphs, making it difficult to predict a priori what the expected behavior of $\mathrm{ex}(G_{n,p}^{r},F)$ should be for a given $r$-graph $F$. Moreover, this lack of a general conjecture for the random Turán problem of hypergraphs turns out to be closely related to the lack of an analog of Sidorenko’s conjecture for hypergraphs. To discuss this further, we need to formally define what it means for a hypergraph to be Sidorenko.
Definition 1.
A homomorphism from an $r$-graph $F$ to an $r$-graph $H$ is a map $\phi:V(F)\to V(H)$ such that $\phi(e)$ is an edge of $H$ whenever $e$ is an edge of $F$. We let $\hom(F,H)$ denote the number of homomorphisms from $F$ to $H$ and define the homomorphism density
$t_{F}(H)=\frac{\hom(F,H)}{v(H)^{v(F)}}.$
We say that an $r$-graph $F$ is Sidorenko if for every $r$-graph $H$ we have
$t_{F}(H)\geq t_{K_{r}^{r}}(H)^{|F|},$
where $K_{r}^{r}$ is the $r$-graph consisting of a single edge.
The study of Sidorenko hypergraphs is an extremely active area of research within extremal combinatorics, especially in the context of graphs [3, 6, 7, 5, 9, 11, 12, 15, 17, 18, 37], with the most outstanding problem in this direction being the following.
Conjecture 2 (Sidorenko’s conjecture [35, 34]).
Every bipartite graph is Sidorenko.
It is well known that Sidorenko’s conjecture does not extend to hypergraphs, in the sense that there exist $r$-partite $r$-graphs $F$ which are not Sidorenko for all $r\geq 3$. For such hypergraphs, it was shown by Conlon, Lee, and Sidorenko [8] that $\mathrm{ex}(n,F)$ is always strictly larger than the trivial bound given by a simple deletion argument. This result was later extend to the random Turán setting in [29], where in particular it was shown that essentially any counterexample to Sidorenko’s conjecture would provide a counterexample to 1, and more generally that any non-Sidorenko $r$-partite $r$-graph would provide a counterexample to the hypergraph analog of 1.
1.3 Hypergraph Expansions
The discussion above illustrates that getting a complete understanding of random Turán numbers of arbitrary hypergraphs is complicated by the existence of non-Sidorenko hypergraphs. To partially overcome this, we will restrict ourselves in this paper to the following class of hypergraphs which is known to behave well with respect to both random Turán numbers and being Sidorenko.
@@ 1.3 Hypergraph Expansions
One nice features about hypergraph expansions is that their Turán numbers are easy to compute for large $r$. In particular, it is straightforward to show for any $r_{0}$-graph $F$ which does not contain a vertex contained in every edge of $F$ that $\mathrm{ex}(n,F^{(r)})=\Theta_{r}(n^{r-1})$ for all $r$ sufficiently large in terms of $F$. In view of this, it is natural to ask if one can also easily establish bounds for random Turán numbers of expansions, at least in the case when $r$ is large and $p$ is relatively close to 1. And indeed, there have been a number of works studying special cases of this problem [25, 30, 26, 27], with perhaps the most well studied case being that of loose even cycles $C_{2\ell}^{(r)}$.
The particular problem of studying loose even cycles was initially studied by Mubayi and Yepremyan [25], with later independent work also being done by Nie [27]. In total, these two streams of research culminated in the following set of tight bounds for this problem whenever the uniformity is at least $4$.
@@
We emphasize here the peculiar situation that tight bounds for loose cycles are unknown when $r=3$, despite tight bounds being known for $r=2$ due to [22] and for $r\geq 4$ due to 1.1. We also note that very recently, Jiang and Longbrake [14] improved upon the bounds of 1.1 by obtaining tight asymptotic bounds for $r\geq 5$ and $p$ sufficiently large through developing a supersaturation version of the celebrated delta-system method.
For general hypergraph expansions, the behavior of $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ when both $r$ and $p$ are large was determined by Nie [26].
Theorem 1.2 (Theorem 1.3, [26]).
If $F$ is an $r_{0}$-graph, then for all $r\geq v(F)$ and
$p\geq n^{-\frac{r-v(F)}{r-v(F)+1}\cdot\frac{1}{d_{r}(F^{(r)})}}(\log n)^{\Omega(1)}$, we have a.a.s.
$\mathrm{ex}(G^{r}_{n,p},F^{(r)})=\Theta\left(pn^{r-1}\right).$
In fact, a somewhat stronger result is proven in [26] where the $v(F)$ term throughout 1.2 can be replaced by a somewhat smaller integer $k$, but in either case the theorem fails to determine the behavior of $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ for all ranges of $p$ for any hypergraph $F$ outside of certain classes of hypergraph trees.
This failure to determine the random Turán number for all values of $p$ and nearly all hypergraphs is largely due to the fact that $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ can exhibit wildly differing behaviors for different choices of $F$. For example, if $F$ is an even cycle, then 1.1 shows that $\mathrm{ex}(G_{n,p}^{r},F^{(r)})=\Theta(pn^{r-1})$ for all $p\gg n^{\frac{-1}{d_{r}(F^{(r)})}}$. On the other hand, if $F=C_{3}$ then it is known [30, 26] that $\mathrm{ex}(G_{n,p},F^{(r)})\gg pn^{r-1}$ for $n^{\frac{-1}{d_{r}(F^{(r)})}}\ll p\ll n^{-\frac{r-v(F)}{r-v(F)+1}\cdot\frac{1}{d_{r}(F^{(r)})}}$, which in particular shows that the range of $p$ for which 1.2 holds can not be increased for arbitrary $F$. More generally, it is known that $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ will be substantially larger than $pn^{r-1}$ for certain ranges of $p\gg n^{\frac{-1}{d_{r}(F^{(r)})}}$ whenever $F^{(r)}$ is not Sidorenko, and in these cases the behavior of $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ is largely unknown even in the simple case of $F=C_{5}$.
2 Main Results
Our main result is a significant extension of 1.2, essentially classifying which hypergraphs $F$ are such that the random Turán numbers $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ for large $r$ exhibits three simple ranges of behavior analogous to the ranges predicted in 1 for graphs. To state this classification, we say that an $r_{0}$-graph $F$ with $|F|\geq 3$ is strictly $r_{0}$-balanced if $d_{r_{0}}(F)=\frac{|F|-1}{v(F)-r_{0}}$ and if $\frac{|F^{\prime}|-1}{v(F^{\prime})-r_{0}}<d_{r_{0}}(F)$ for every proper subgraph $F^{\prime}\subseteq F$ with at least two edges. We also let $\Delta(F)$ denote the maximum degree of $F$.
@@ 2 Main Results
Let $F$ be a strictly $r_{0}$-balanced $r_{0}$-graph with $2\leq\Delta(F)<|F|$. If $F$ is Sidorenko, then for all $r>|F|^{2}v(F)r_{0}$, we have a.a.s.
$\mathrm{ex}(G^{r}_{n,p},F^{(r)})=\left\{\begin{aligned} &\max\left\{\Theta(pn^{r-1}),n^{r_{0}-\frac{1}{d_{r_{0}}(F)}}(\log n)^{\Theta(1)}\right\},~{}~{}&\text{if}~{}p\gg n^{-r+r_{0}-\frac{1}{d_{r_{0}}(F)}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-r+r_{0}-\frac{1}{d_{r_{0}}(F)}}.\\ \end{aligned}\right.$
Moreover, if there does not exist any $k$ such that $F^{(k)}$ is Sidorenko, then these bounds fail to hold for all $r$.
As an aside, it is unknown whether or not there exist hypergraphs $F$ which are not Sidorenko and which have some expansion $F^{(k)}$ that is Sidorenko [29, Question 5.2], so the hypothesis in the“moreover” part of this statement may in fact be equivalent to saying that $F$ is not Sidorenko.
While the bound $r>|F|^{2}v(F)r_{0}$ from 2.1 is not very large given the high level of generality of the theorem statement, it is natural to ask if these bounds can be improved for specific choices of $F$. In particular, the following problem is of interest.
Question 1.
Which Sidorenko $r_{0}$-graphs $F$ are such that $\mathrm{ex}(n,F)=\omega(n^{r_{0}-1})$ and are such that the tight bounds of 2.1 hold for every $r$ with $\mathrm{ex}(n,F^{(r)})=\Theta(n^{r-1})$?
The hypothesis that $\mathrm{ex}(n,F^{(r)})=\Theta(n^{r-1})$ is necessary for the bounds of 2.1 to hold, as otherwise the stated bound for $p=1$ would be incorrect. The condition $\mathrm{ex}(n,F)=\omega(n^{r_{0}-1})$ is introduced primarily to avoid trivial solutions to this problem, since without this condition, every large expansion of a Sidorenko hypergraph would give a positive answer by 2.1.
As far as we are aware, there are no known examples for which either positive or negative answers to 1 are known, and in particular, it may very well be the case that 1 has a positive answer for every Sidorenko hypergraph. The closest the current literature comes to verifying any case of 1 is 1.1 for $F$ an even cycle, giving tight bounds for all $r\geq 4$ (c.f. 1 which asks for $r\geq 3$ in this case).
Although we remain unable to answer 1 even for the simplest even cycle $C_{4}=K_{2,2}$, we can give a positive answer for every other $K_{2,t}$, giving the first (infinite number of) positive answers to 1.
Theorem 2.2.
For all integers $r,t\geq 3$ and all $0<p=p(n)\leq 1$ we have a.a.s.
@@
More generally, we can significantly improve upon the dependency on $r$ from 2.1 for all complete bipartite graphs $K_{s,t}$, though optimal bounds only hold for $s=2$.
@@
$\mathrm{ex}(G^{r}_{n,p},K_{s,t}^{(r)})=\left\{\begin{aligned} &\max\{\Theta(pn^{r-1}),n^{2-\frac{s+t-2}{st-1}}(\log n)^{\Theta(1)}\},~{}~{}&\text{if}~{}p\gg n^{-r+2-\frac{s+t-2}{st-1}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-r+2-\frac{s+t-2}{st-1}}.\\ \end{aligned}\right.$
It is possible that these tight bounds for $\mathrm{ex}(G_{n,p}^{r},K_{s,t}^{(r)})$ continue to hold for any $r$ such that (${r\choose 2}\geq s$, as this is precisely the range where $\mathrm{ex}(n,K_{s,t}^{(r)})=\Theta(n^{r-1})$ for $t$ sufficiently large [19, Theorem 1.4], but 2.3 is the best bound for $r$ that can be obtained with our general methods and the known (tight) random Turán bounds for the graph $K_{s,t}$ [22]. This being said, in forthcoming work [28], we use a significantly more technical argument tailored specifically towards $K_{s,t}$ expansions in order to go beyond 2.3 and prove the same result for $r$ at least roughly $12s/3+$ 1, in particular giving another positive answer to 1 for $K_{4,t}$ and $t$ sufficiently large.
Finally, we give a positive solution to 1 for theta graphs with sufficiently many paths. Here the theta graph $\theta_{a,b}$ is defined to be the graph consisting of $a$ internally disjoint paths of length $b$ between a common pair of vertices $u,v$. For example, $\theta_{2,b}=C_{2b}$ and $\theta_{a,2}=K_{2,a}$. Our general methods give good bounds on $\theta_{a,b}^{(r)}$ whenever effective bounds for the random Turán number of $\theta_{a,b}$ are known and $a\geq 3$ (see 6.7), and in particular, recent results of McKinley and Spiro [20] give the following.
@@
2.1 General Tools for Random Turán Problems
In the process of proving the results detailed above, we will establish a number of general tools for random Turán problems (listed out below) which we expect will be of use in future study into this area, especially for problems related to hypergraph expansions. In particular, we note that this collection of results are enough to fairly easily recover every previously known upper bound for random Turán problems of expansions [25, 30, 26, 27].
We begin with a very basic result that esatablishes effective bounds for arbitrary hypergraphs $F$.
Proposition 2.5.
$\mathrm{ex}(G_{n,p}^{r},F)=(1+o(1))p{n\choose r}.$
If $p\gg n^{-\frac{1}{d_{r}(F)}}$, then a.a.s.
With 2.5 in hand, all that is needed to prove tight bounds for $\mathrm{ex}(G_{n,p}^{r},F)$ are effective upper bounds on the problem whenever $p\gg n^{-1/d_{r}(F)}$, which is usually the most challenging component of random Turán problems. Establishing such bounds typically requires proving a “balanced supersaturation” result, which informally says that if an $r$-graph $H$ has many edges, then one can find a collection $\mathcal{H}$ of copies of $F$ in $H$ which is “spread out” in the sense that no set of edges in $H$ lies in too many copies of $F$ in $\mathcal{H}$.
More precisely, we will make use of the following definition, where here for a hypergraph $\mathcal{H}$ and $S\subseteq V(\mathcal{H})$ we define the degree $\deg_{\mathcal{H}}(S)$ to be the number of edges of $\mathcal{H}$ containing $S$, and we define $\Delta_{i}(\mathcal{H})=\max\{\deg_{\mathcal{H}}(S):S\subseteq V(\mathcal{H}),\ |S|=i\}$.